Ali’s budget line relating good X and good Y has intercept of 50 unit of good X and 20 units of good Y. if the price of good X is 12, what is Ali’s income? What is the price of good Y? What is then slope of budget line?


1
Expert's answer
2021-04-26T19:20:52-0400

Budget line of good X &Y has intercept then individual A income is

"Q_X=50" units, "Q_Y=20" units, "P_X="$"12"

"Income=P_X\times Q_X+P_Y \times Q_Y"

When the budget line has intercept "Q_X=50" units and "P_X=" $"12"

then, income will be,

"=50\times 12"

"income=" $"600"


The price (p) of good Y is the point of interception with X-axis

"600=12\times 0+P_Y \times 20"

"P_Y=" $"30"

The slope of budget is

"600=12\times Q_X+30\times Q_Y"

"30Q_Y=600-15Q_X"

"Q_Y=20-0.5Q_X"

slope of the budget line"=-0.5"